# EPOCHS IV: SED Modelling Assumptions and their impact on the Stellar Mass Function at using PEARLS and public JWST observations

### v1

@@ 7 Conclusions
In this paper we present an investigation into the properties of the EPOCHS v1 sample of 1120 high-redshift galaxies at 6.5 taken from a uniform reduction of 187 arcmin of JWST data, including the GTO program PEARLS as well as other public ERS/GO JWST programs.
@@
4. Across all of the fitted models, the highest mass galaxies in our sample are ‘Little Red Dots’, with inferred masses of 10 M at . These galaxies dominate the highest mass bins of our galaxy stellar mass function (GSMF) estimates, so understanding their true stellar populations and accounting for the likely contribution of AGN (Greene et al., 2023; Furtak et al., 2023, 2023) will be essential to more accurately constrain further GSMF estimates.
@@
This is only the beginning of GSMF estimates at , and the use of ultra-deep observations (the second NGDEEP epoch (Bagley et al., 2023a), the JADES Origins Field (Eisenstein et al., 2023b; Robertson et al., 2023a) and others) and magnification by lensing clusters (PEARLS, UNCOVER, CANUCS; Windhorst et al., 2023; Bezanson et al., 2022; Willott et al., 2023b) will help constrain the GSMF at stellar masses below our completeness limit of M, whilst widefield surveys (e.g. PRIMER, UNCOVER, Cosmos-Webb; Dunlop et al., 2021; Bezanson et al., 2022; Kartaltepe et al., 2021) will add area and rare higher-mass sources. Deep MIRI F560W or F770W observations (e.g. the MIRI HUDF survey, Norgaard-Nielsen & Perez-Gonzalez, 2017) will be crucial to provide better constraints on stellar mass estimates at these redshifts by extending the wavelength range further into the rest-frame optical, although the sensitivity of MIRI decreases rapidly with increasing wavelength (e.g. Wang et al., 2024). More complete NIRSpec coverage is also important to identify interlopers, confirm photometric redshifts and distinguish between AGN emission and star forming galaxies.
@@
All of the raw JWST data used in this work are the same as used in Adams et al. (2023) and can be accessed via this MAST DOI: DOI 10.17909/5h64-g193. All proprietary data from the PEARLS program will all become accessible over 2024. Catalogues for all high- galaxies will be published with the EPOCHS I paper (Conselice et. al., in prep). The fiducial GSMF and SMD results from this work are availale on GitHub, and results for our alternative models will be made available upon request.
Appendix A z = 7 Stellar Mass Function without ‘Little Red Dots’
@@
We exclude these 34 candidates from our sample and reconstruct the stellar mass function at . No other changes are made to our GSMF construction or fitting procedures. Figure 14 shows the GSMF derived without including any ‘Little Red Dots’. Compared to the fiducial GSMF, this removes the two highest mass bins entirely, which demonstrates our reliance on these galaxies at the high mass regime. In terms of the derived Schechter parameters, the exponential mass cutoff M, which is not well constrained, decreases from 11.57 to 10.64 when we exclude the LRDs. The median posterior and for the GSMF without LRDs are and respectively.
Appendix B Tabulated Schechter parameters for alternative GSMF estimates at z = 10.5
@@
| SED Fitting Tool | SFH Model | IMF | M | ||
@@
Appendix C z = 12.5 GSMF with no contamination limit
Our fiducial GSMF applies a 50% contamination limit on all galaxies. Given the fields a galaxy is selected in and its stellar mass, our JAGUAR contamination simulation computes a likelihood of contamination, based on simulated galaxies with the same stellar mass. The highest contamination is seen in the GSMF, for the M/M bin, and results in several galaxies being removed from our fiducial GSMF estimate. As the predictions of JAGUAR are uncertain at these redshifts, it is hard to judge how accurate our contamination predictions are. In Figure 15 we have recomputed the GSMF with no contamination limit, which boosts the lower stellar mass bin by 0.3 dex. This brings it closer to the predictions of Li et al. (2023)’s Feedback Free Model (FFB), which has higher star formation efficiency than most models. The FFB model shown is for specifically, and the models with higher SFE () overpredict the GSMF at this redshift compared to our observations. If our contamination is over-estimated in this redshift bin, then the FFB model of Li et al. (2023) or Mauerhofer & Dayal (2023)’s DELPHI model provide the closest predictions, suggesting high but not extreme star formation efficiency is required to produce the observed GSMF at this redshift.

### v3

@@ 7 Conclusions
In this paper we present an investigation into the properties of the EPOCHS v1 sample of 1120 high-redshift galaxies at 6.5 taken from a uniform reduction of 187 arcmin2 of JWST data, including the GTO program PEARLS as well as other public ERS/GO JWST programs.
@@
4. Across all of the fitted models, the highest mass galaxies in our sample are ‘Little Red Dots’, with inferred masses of 1010 M⊙ at . These galaxies dominate the highest mass bins of our galaxy stellar mass function (GSMF) estimates, so understanding their true stellar populations and accounting for the likely contribution of AGN (Greene et al., 2023; Furtak et al., 2023, 2023; Williams et al., 2024) will be essential to more accurately constrain further GSMF estimates.
@@
This is only the beginning of GSMF estimates at , and the use of ultra-deep observations (the second NGDEEP epoch (Bagley et al., 2023a), the JADES Origins Field (Eisenstein et al., 2023b; Robertson et al., 2023a) and others) and magnification by lensing clusters (PEARLS, UNCOVER, CANUCS; Windhorst et al., 2023; Bezanson et al., 2022; Willott et al., 2023b) will help constrain the GSMF at stellar masses below our completeness limit of M⊙, whilst widefield surveys (e.g. PRIMER, UNCOVER, Cosmos-Webb; Dunlop et al., 2021; Bezanson et al., 2022; Kartaltepe et al., 2021) will add area and rare higher-mass sources. Deep MIRI F560W or F770W observations (e.g. the MIRI HUDF survey, Norgaard-Nielsen & Perez-Gonzalez, 2017) will be crucial to provide better constraints on stellar mass estimates at these redshifts by extending the wavelength range further into the rest-frame optical, although the sensitivity of MIRI decreases rapidly with increasing wavelength (e.g. Wang et al., 2024). More complete NIRSpec coverage is also important to identify interlopers, confirm photometric redshifts and distinguish between AGN emission and star forming galaxies.
@@
All of the raw JWST data used in this work are the same as used in Adams et al. (2023) and can be accessed via this MAST DOI: DOI 10.17909/5h64-g193. All proprietary data from the PEARLS program will all become accessible over 2024. Catalogues for all high- galaxies will be published with the EPOCHS I paper (Conselice et al., 2024). The fiducial GSMF and SMD results from this work are availale on GitHub, and results for our alternative models will be made available upon request.
Appendix A Effect of SED Fitting Assumptions on Derived SPS Quantities
Here we give a more detailed comparison between our fiducial Bagpipes results and the alternative models discussed in § 4.1. In Figure 13 we show the equivalent of Figure 1 for other derived parameters, in order to understand what drives the observed discrepancies in stellar mass between different models.
A.1 Impact of Priors
The first comparison we make is a substitution of the default logarithmic prior on the V-band dust extinction, A, to a uniform prior. A uniform prior favors higher dust extinctions, and the largest difference to the fiducial model is seen in galaxies with A mag, where the dust content is poorly constrained. We see good agreement in photo-, along with a few extreme outliers, which is unsurprising given the informative redshift prior from EAZY-py we use. The stellar mass offset is shown in light blue in the top plot of Figure 1 and shows good agreement within the posterior uncertainties. Galaxies with higher levels of dust are found to be significantly younger and more star-forming with a uniform dust prior, as can been see in Figure 13 suggesting the inferred star formation histories are dependent on the dust prior. Figure 13 also shows that the measured dust extinction which is influenced by the prior is primarily at , with both models agreeing on the dust extinction for the galaxies with higher dust extinction. Comparison of the best-fitting values shows that both models are equally well-fitted to the photometry.
We also test the impact of our metallicity prior, which is logarithmic in our fiducial Bagpipes model. This favours low metallicity, which we expect in the early Universe. Here we exchange this prior for a uniform distribution which favours higher metallicity, shown in orange in the top plot of Figure 1. We see little overall impact on the stellar mass from the metallicity prior, with individual galaxies scattering up to 0.5 dex and the majority consistent with results from our fiducial Bagpipes run. The metallicity itself, which is extremely difficult to constrain from photometry, is highly influenced by the prior chosen, as can be seen in Figure 13, although this doesn’t appear to systematically impact the star formation history or dust extinction.
A.2 The Assumed Dust Law
Our fiducial Bagpipes model assumes a simple one-component Calzetti et al. (2000) dust law. The slope of the dust law is known to vary in some galaxies, and numerous alternative models (e.g. Charlot & Fall, 2000; Salim et al., 2018) including additional parameters have been suggested. Charlot & Fall (2000) fit a two-component dust law, with different amounts of extinction for young ( Myr) and old stellar populations, to account for dust in stellar birth clouds. The resultant deviations in stellar mass are shown in magenta in the top plot of Figure 1, with a significant deviation at the highest stellar masses, which can be seen in the LOWESS trend. For galaxies with a fiducial stellar mass of M⊙, the majority are found to have 1 dex larger stellar masses with the Charlot & Fall (2000) dust law. However for these galaxies the goodness of fit is considerably poorer, with with the Charlot & Fall (2000) model.
Despite the same prior, the dust attenuation (A) posteriors are quite different between the two models, due to the degeneracy between A and slope, as can be seen in Figure 13. Galaxies with moderate dust extinction in the fiducial results ( mag) are typically found to have very little dust extinction on the old stellar population, with higher dust extinction on the young stellar populations. The majority of galaxies favor a steeper slope () for the attenuation power-law than given by Calzetti et al. (2000), in this mode with a distribution centered on n, which is 2 from the prior value of 0.70.
Salim et al. (2018) allow a deviation in slope compared to the Calzetti et al. (2000) and an additional UV bump at . The UV bump in the Salim et al. (2018) model is driven primarily by Polycyclic Aromatic Hydrocarbon (PAH) emission, which is expected when emission is strongly reprocessed. We replace the Calzetti et al. (2000) dust law with the Salim et al. (2018) model, and fit for these additional components, as shown in § 4.1. We find that the inferred stellar mass can increase by 1 dex in some cases, particularly at the highest fiducial stellar mass, but the average offset (see the LOWESS fit in Figure 1 is smaller than for the Charlot & Fall (2000) case. In a number of cases the goodness of fit for the Salim et al. (2018) model is considerably better than our fiducial model, with in a number of galaxies.
A.3 Comparison to ‘delayed’ SFH
To test the consistency of galaxy stellar mass estimates with different parametric star formation histories, we replace our ‘lognormal’ SFH with another commonly used SFH; a delayed- SFH. We find systematically slightly lower stellar masses above 108 M⊙, but reasonable agreement at the lowest stellar masses. This is likely due to the log-uniform prior on age used in the delayed- model, which results in a systematically younger stellar population, as can be seen in Figure 13. We also see some impact on the dust attenuation, likely driven by the change in inferred SFH, with systematically higher dust attenuation inferred with this SFH. Comparison of the goodness of fit via the parameter suggests the models typically have slightly poorer fits than the fiducial model, but in some cases the stellar masses are reduced by 0.5 dex with very little impact on the goodness of fit.
A.4 Comparison to “continuity bursty” SFH
Here we replace our fiducial Bagpipes ‘lognormal’ SFH with a “continuity bursty” SFH model described in detail in § 4.1. We reproduce the result of Tacchella et al. (2022) that this SFH increases the galaxy stellar mass, finding an average increase of 0.2 dex. The stellar mass discrepancy between the two models is shown in purple in the lower figure of 1. In individual cases the increase in stellar mass can reach 1 dex, with only small changes in photo-. The largest offsets in stellar mass are typically seen for galaxies with a fiducial stellar mass of M⊙, with the highest fiducial stellar mass galaxies ( M⊙) seeing considerably smaller increases.
Looking at the other derived properties in Figure 13, we also see significant scatter in dust extinction, where the galaxies with highest extinctions in our fiducial model found to contain significantly less dust with this SFH model. For galaxies older than 50 Myr in our fiducial model, they are typically found to be systematically younger with this non-parametric SFH, but some of the youngest galaxies in the fiducial model are found to be much older.
A.5 Comparison to BPASS SPS Models
By default Bagpipes uses the 2016 version of Bruzual & Charlot (2003)’s (BC03) stellar population synthesis (SPS) models. However Bagpipes can also employ Binary Population and Spectral Synthesis (BPASS, Stanway & Eldridge, 2018) models. Specifically we use models generated with v2.2.1 of BPASS, with the default IMF (slope of 1.35, 300 M⋆). The IMF parametrization of this SPS model differs slightly from the Kroupa (2001) IMF used in the BC03 models, causing an intrinsic offset in stellar mass. However as shown in orange in the lower plot of Figure 1, the comparison of the mass estimates is more complex. At the lowest and highest stellar masses ( and ), we see systematically lower stellar masses than the fiducial Bagpipes model, which is what we would expect based the IMF difference alone. However for the majority of the sample, which falls between these two mass extremes, we see a significantly larger offset in stellar masses, which increases with redshift, as shown by the LOWESS trend.
In Figure 13 we also see the significant impact the BPASS SPS model has on the derived dust extinction, age, and metallicity. Despite the same prior and dust model, when the BPASS SPS model is used the dust extinction AV is systematically much lower, often inferring essentially no dust extinction. Galaxy ages are also found to be systematically much larger, often by 100 Myr.
A.6 Other Comparisons
2. “continuity” non-parametric SFH model’: This variation tests the other non-parametric SFH model introduced in Leja et al. (2019) and Tacchella et al. (2022), which more tightly constrains the SFR in neighbouring time bins to force more smoothly varying star formation histories than the “continuity bursty” model. We find overall similar behaviour to the “continuity bursty” model, with systematically higher masses and higher mass-weighted age. The models typically have worse statistics than the “continuity bursty” model fits potentially indicating that more stochastic star formation histories are preferred for the majority of galaxies, as suggested in the literature (Faucher-Giguère, 2018; Looser et al., 2023; Asada et al., 2023; Endsley et al., 2023b).
A.7 Comparison to Prospector
For the comparison of the parametric SFH models (delayed- for Prospector, lognormal for Bagpipes), we see systematically larger photo- and stellar mass estimates with Prospector. Whilst in the Prospector model we allow the IGM attenuation to vary, which could impact photo- estimates, we do not see the same offset with the non-parametric SFH model, where the IGM attenuation is also allowed to vary. Both SED-fitting tools are given the same redshift prior from EAZY-py. For both example galaxy SEDs in Figure 3 that Prospector instead prefers higher- solutions than Bagpipes because it is inferring the presence of Lyman- emission. Whilst Lyman- emitters have been found at (e.g. Bunker et al., 2023b), we do not expect to observe Lyman- from the majority of galaxies at these redshifts due to the attenuation from neutral hydrogen during the Epoch of Reionization, and a systematic photo- offset from Prospector is likely for galaxies without Lyman- emission. However given that we see a photo offset primarily for the ’parametric SFH’ only, Lyman- emission is unlikely to be the only cause of the offset as the Prospector SEDs in Figure 3 show Lyman- emission for both SFH models. This offset will generally cause a slight increase in stellar mass estimates, as a more distant galaxy must be intrinsically brighter. We see a median increase of 0.43 dex in stellar mass.
Appendix B z = 7 Stellar Mass Function without ‘Little Red Dots’
@@ A.7 Comparison to Prospector
We exclude these 34 candidates from our sample and reconstruct the stellar mass function at . No other changes are made to our GSMF construction or fitting procedures. Figure 15 shows the GSMF derived without including any ‘Little Red Dots’, following the cut of Kokorev et al. (2024). Compared to the fiducial GSMF, this removes the two highest mass bins entirely, which demonstrates our reliance on these galaxies at the high mass regime. In terms of the derived Schechter parameters, the exponential mass cutoff M⋆, which is not well constrained, decreases from 11.57 to 10.64 when we exclude the LRDs. The median posterior and for the GSMF without LRDs are and respectively. We also test the more restrictive color cut of Williams et al. (2024), which require a stronger F277W-F444W color criteria in order to provide an 80% AGN purity (Greene et al., 2023). The majority of the LRDs we find above do not meet this criteria, with only 8/34 having F277W-F444W 1.6 mag, but these LRDs with the reddest colors also typically have the highest inferred stellar masses, so excluding just these 8 objects has a noticeable impact on the GSMF, as shown in Figure 15.
Appendix C Tabulated Schechter parameters for alternative GSMF estimates at z = 10.5
@@
SED Fitting Tool | SFH Model | IMF | M⋆ | ||
@@
Appendix D z = 12.5 GSMF with no contamination limit
Our fiducial GSMF applies a 50% contamination limit on all galaxies. Given the fields a galaxy is selected in and its stellar mass, our JAGUAR contamination simulation computes a likelihood of contamination, based on simulated galaxies with the same stellar mass. The highest contamination is seen in the GSMF, for the M⋆/M bin, and results in several galaxies being removed from our fiducial GSMF estimate. As the predictions of JAGUAR are uncertain at these redshifts, it is hard to judge how accurate our contamination predictions are. In Figure 16 we have recomputed the GSMF with no contamination limit, which boosts the lower stellar mass bin by 0.3 dex. This brings it closer to the predictions of Li et al. (2023)’s Feedback Free Model (FFB), which has higher star formation efficiency than most models. The FFB model shown is for specifically, and the models with higher SFE () overpredict the GSMF at this redshift compared to our observations. If our contamination is over-estimated in this redshift bin, then the FFB model of Li et al. (2023) or Mauerhofer & Dayal (2023)’s DELPHI model provide the closest predictions, suggesting high but not extreme star formation efficiency is required to produce the observed GSMF at this redshift.